SPECTROSCOPY AND DISSOCIATION DYNAMICS OF THE 1-METHYLALLYL RADICAL

M. GASSER, J. A. FREY, J. M. HOSTETTLER and \underline{A. BACH}, Laboratorium für Organische Chemie, ETH Zürich, CH-8093, Switzerland.

The Ã ← X̃ band system of jet-cooled 1-methylallyl radical (C_{4}H_{7}) was observed for the first time using resonance-enhanced multiphoton ionization combined with electronic ground state depletion spectroscopy. Analysis of the vibronic structure reveals transitions to the nonplanar Ã valence excited state with an electronic origin for the Z-isomer of 1-methylallyl at 23 979 cm^{−1}. Time- and frequency-resolved photoionization of the hydrogen atom product from electronically excited 1-methylallyl radical and its isotopologue CH_{3}C_{3}D_{4} provides information on the dissociation dynamics. The measured dissociation rates and kinetic energy release combined with results from high level ab initio calculations suggests unimolecular decompostion to 1,3-butadiene and hydrogen with no evidence for nonstatistical behavior in dissociation.